OpenStudy (anonymous):

A pineapple was found to have two sets of opposing spirals formed by its leaves. The clockwise spiral has x leaves. The number of leaves in the counterclockwise direction is 13 more than that in the clockwise direction. These leaf numbers are consecutive numbers in the Fibonacci sequence. Which equation is true for the number of leaves in the two spirals?

OpenStudy (anonymous):

You didn't give any options, but the number of spirals has to be 21 and 34. 13 more means that you added 13. The only time you add 13 in the fibonacci sequence is 21 and 13 to get 34 (not 8 and 13 to get 21 since 8 and 21 are nonconsecutive). You didn't give any options for your question though so I just solved for x.

OpenStudy (anonymous):

A pineapple was found to have two sets of opposing spirals formed by its leaves. The clockwise spiral has x leaves. The number of leaves in the counterclockwise direction is 5 more than that in the clockwise direction. These leaf numbers are consecutive numbers in the Fibonacci sequence. which equation is true for the number of leaves in the two spirals? x+5/5=1.6 x/5=1.6 5/x=1.6 x+5/x=1.6